OpenStudy (anonymous):

Simplify the following expression. √-24 √-81

OpenStudy (anonymous):

square root of -24 multiplied by the square root of -81

OpenStudy (anonymous):

no it says simplify the two

OpenStudy (anonymous):

oh ok misread sorry

OpenStudy (anonymous):

think it has to be in terms of i

OpenStudy (anonymous):

whats the square root of 24?

OpenStudy (anonymous):

is it 6*4=24

Directrix (directrix):

@artishaalston √-24 √-81 = These two numbers are not real. The first step is to write √-24 as √-24 * √-1 and √-81 as √-1 * √81. √-24 √-81 = √-24 * √-1* √-1 * √81 = What is √-1* √-1 when simplified?

OpenStudy (anonymous):

no its a number multiplied by its self

OpenStudy (anonymous):

nothing time it selfs gives you 24

OpenStudy (anonymous):

i